OpenStudy (anonymous):

Body centered cubic lattice question - A metallic iron has a body centered lattice. The density of iron is 7.87 g/cm3, what is the edge length of the unit cell of this sample of iron? Please explain and show work! Thank you!

OpenStudy (chmvijay):

ρ =nA / VN n=2 for BCC A = atomic eight of Fe Volume of V =( 4/R suare root of 3 )^3 Density = 7.87 g/cm3 N is Avogadro number find R from the above equation

OpenStudy (chmvijay):

then the valu eof R u put in this equation to get ur result that is a=? a = 4/\[\sqrt{3R}\]